the average number of pages in a novel is 326 with a standard deviation of 24 pages. If a sample of 50 novels is randomly chosen, what is the probability that the average number of pages in these books is between 319 and 331?





1
Expert's answer
2022-04-05T07:29:18-0400

"P(319<X<331)=P(\frac{319-326}{24}<Z<\frac{331-326}{24})=P(-0.29<Z<0.21)"

P(319<X<331)=0.58317-0.38591=0.19726